# ChainBuddy: An AI Agent System for Generating LLM Pipelines

### v1

@@ 5.2. Qualitative results
We first expand on these findings in the sections below and then discuss the interaction patterns identified by the study
@@ 5.2.5. Accelerating workflow
Participants remarked that the assistant significantly accelerated their workflow. P11 remarked that ”with the assistant, it’s really quick. It gives you the ideas, and it makes all the necessary connections on the platform, so you save up that manual time of creating the whole workflow. And on top of that, it also saves mental efforts to start with the prompt.” Similarly, P9 felt that while there wasn’t a significant difference in their success rate and solution, ”it was just faster with ChainBuddy”.
@@ 5.2.10. Use cases in free exploration time
During the free exploration time, participants explored a diverse range of ideas, showcasing creative and varied uses in collaborating with the assistant to address specific needs and interests. The diverse range of topics reflected a similar diverse range in the study of ChainForge (Arawjo et al., 2024), the baseline interface. Participant usage fell into three major categories: comparing and exploring different prompts, evaluating behavior across different LLMs, and testing LLMs for bias and handling of sensitive topics. The full list of participant ideas is listed in Appendix A.
@@ 5.3. Limitations
Our within-subjects usability study had a small sample size and compared to a single baseline. It could be that for a different LLM pipeline interface, we might have seen different results. We did not restrict users by prior knowledge of the baseline interface—it could be that expert ChainForge users would feel differently about the assistant’s capabilities. For various constraints common to usability studies, we also only examined two tasks that both involved comparing across prompts, even though ChainBuddy is capable of more types of tasks and LLM pipelines, such as comparing responses across models or setting up data processing workflows. An alternate study design might explicitly try three or more very different tasks with a larger sample size, to better understand when ChainBuddy is and is not useful. We also relied on rough indicators of performance; a more robust evaluation might have a team of external raters grade the quality of created flows on a scale. Finally, our quantitative findings should also be taken with a grain of salt: the complex interaction effects we observed for Ease of Use and Successfulness in particular may be due to random noise. Nevertheless, we believe that the body of evidence is clear, in that all tests of significance leaned in favor of ChainBuddy, and qualitative feedback was overwhelmingly positive. We intend to run a follow-up study post-deployment, investigating a broader range of tasks and real-world use cases.
@@ 6. Discussion
One interesting null result is time. We did not find significant qualitative evidence that ChainBuddy helped people solve the task faster, as might be expected (although participants did express this feeling in post-interviews). However, three things to note. First, ChainBuddy takes time to generate flows, on the order of 10 seconds or greater. Second, our times are subjective, with a participant-reported indication of being “done.” Participants using the assistant spent time revising, running or inspecting the generated flows, or clarifying their intent with the assistant. Third, qualitative data indicates that some participants felt they could go further in their analysis with the assistant, compared to without. So, though participants spent around the same time per condition, what they spent that time on differed.
More broadly than LLM pipelines, our work contributes to a growing body of literature on AI agents integrated into software platforms to assist users in implementing ideas (e.g., an agent that creates Powerpoint slides, visualizations from datasets, or dynamic widgets (Vaithilingam et al., 2024b; Beasley and Abouzied, 2024; Zheng et al., 2022)). Our study provides evidence for the intuitive hypothesis that users perceive that AI agents improve workload by reducing the mental and physical burdens placed on them when interacting with software—clicking and dragging, planning what to do. However, our quantitative findings indicate that for the most part, users perceived they were just as successful without the assistant—even when performance analysis suggests overall performance plummeted without the assistants’ help. This finding resembles the Dunning–Kruger effect (Kruger and Dunning, 1999)—that people who lack the skill to evaluate the quality of their work in a domain may over-estimate its quality—and, combined with worries about being biased by the assistant (discussed below), contributes to growing concerns about non-experts’ over-reliance on AI systems (Buçinca et al., 2021; Bansal et al., 2021). Future studies on software-integrated AI assistants should complement subjective measures with more objective evaluations of user performance (e.g., by external raters or as here, by quantifiable metrics that serve as indicators of performance).
@@ 6.1. Risk of over-reliance on AI assistance for LLM pipeline creation
The reader might ask whether and to what extent the user over-relies on, or is biased by, the assistant’s solutions. Even if the assistant helps the user learn the platform (which some participants indicated), is there a risk of it proposing solutions that are subpar or misleading in some way, outside of more straightforward usability errors? For instance, imagine a user asks ChainBuddy to set up a flow to evaluate LLMs for gender bias. In this task, ChainBuddy usually sets up a flow that compares whether the presence of a social identity marker (male, female, non-binary) changes the LLM success rate on solving a math problem—a common “persona” approach used in benchmarking papers (Cheng et al., 2023). While one way to evaluate gender bias, it certainly does not exhaust all possibilities, nor does the current assistant help the user “scale up” their analysis, choose one solution from alternatives and weigh trade-offs (a learning technique from variation theory (Ling Lo, 2012)), or help them hedge against misinterpretation of random noise. In our study, we did see one error case (Fig. 5) where ChainBuddy constructed a wrong solution that did not compare two prompts. The participant was evidently not able to identify and fix the AI’s mistakes. This raises the question of how future assistants can help the user learn the interface, while mitigating the potential for bias.
@@ 6.2. Future Work
Appendix A Participant Ideas in Free Exploration

### v2

@@ 5.2. Qualitative results
We first expand on these findings in the sections below and then discuss the interaction patterns identified by our user study.
@@ 5.2.4. Reducing effort
[…]
This figure contains two time stack plots comparing participant actions while completing tasks in two conditions: ”Assistant Condition” on the left and ”Control Condition” on the right. Each row represents a participant (P1 to P12), and the x-axis represents the time spent on various actions (measured in minutes from the start of the task). The actions are color-coded in the legend, Light Blue: Chatting with assistant: Represents the time participants spent interacting with the AI assistant to complete tasks. Dark Blue: Waiting for workflow generation. Indicates the waiting time while the assistant generated a workflow. Green: Check generated pipeline. Shows the time spent reviewing the pipeline generated by the assistant. Pink: Edit input data. Represents time spent editing the data inputs used in the task. Red: Generate data from prompt. Captures time spent generating data using manual prompts or the assistant. Yellow-Orange: Edit prompt. Reflects time spent modifying prompts used to perform the task. Purple: Query models and inspect. Represents time querying the model outputs and inspecting results. Dark Purple: Edit Evaluator. Captures the time spent editing evaluation parameters or metrics. Yellow: Run Evaluator and inspect. Represents time running evaluators and inspecting their outputs. Brown: Plot data on graph. Indicates time spent visualizing data or results in graphical form. In the ”Assistant Condition,” participants spend significant time ”Chatting with assistant” and minimal time ”Editing input data.” In the ”Control Condition,” participants spend more time editing inputs and prompts. The figure illustrates a shift in the distribution of actions between the two conditions, with the ”Assistant Condition” reducing manual input tasks.
Figure 6. Interaction traces of participant actions when completing tasks, comparing conditions. Notice the relative dearth of time spent editing input data in the Assistant condition, compared to the Control.
@@ 5.2.5. Accelerating workflow
Participants felt that the assistant significantly accelerated their workflow. P11 remarked that ”with the assistant, it’s really quick. It gives you the ideas, and it makes all the necessary connections on the platform, so you save up that manual time of creating the whole workflow. And on top of that, it also saves mental efforts to start with the prompt.” Similarly, P9 felt that while there wasn’t a significant difference in their success rate and solution, ”it was just faster with ChainBuddy.” Figure 6 shows interaction traces for user actions across both conditions, plotted by time. Most notable is how much effort users in the Control condition spent on preparing and editing input data (TextFields nodes), compared to the Assistant, where only one participant (P9) edited the input data generated by ChainBuddy. Notable too is that half the participants ran evaluators and inspected results when using the Assistant, compared to only one in the Control. Yet, we also see the time users spent chatting with the assistant could sometimes take significant time—in four instances here, it neared or exceeded five minutes.
@@ 5.2.10. Use cases in free exploration time
During the free exploration time, participants explored a diverse range of ideas, showcasing creative and varied uses in collaborating with the assistant to address specific needs and interests. The diverse range of topics reflected a similar diverse range in the study of ChainForge (Arawjo et al., 2024), the baseline interface. Participant usage fell into three major categories: comparing and exploring different prompts, evaluating behavior across different LLMs, and testing LLMs for bias and handling of sensitive topics. The full list of participant ideas is listed in Appendix C.
@@ 5.3. Limitations
Our within-subjects usability study had a small sample size and compared to a single baseline. It could be that for a different LLM pipeline interface, we might have seen different results. We did not restrict users by prior knowledge of the baseline interface—it could be that expert ChainForge users would feel differently about the assistant’s capabilities. For various constraints common to usability studies, we also only examined two tasks that both involved comparing across prompts, even though ChainBuddy is capable of more types of tasks and LLM pipelines, such as comparing responses across models or setting up data processing workflows. An alternate study design might explicitly try three or more very different tasks with a larger sample size, to better understand when ChainBuddy is and is not useful. Finally, our quantitative findings should also be taken with a grain of salt: the complex interaction effects we observed for Ease of Use and Successfulness in particular may be due to random noise. Nevertheless, we believe that the body of evidence is clear, in that all tests of significance leaned in favor of ChainBuddy, and qualitative feedback was overwhelmingly positive. We intend to run a follow-up study post-deployment, investigating a broader range of tasks and real-world use cases.
@@ 6. Discussion
One interesting null result is time. We did not find significant quantitative evidence that ChainBuddy helped people solve the task faster, as might be expected (although participants did express this feeling in post-interviews). However, four things to note. First, ChainBuddy takes time to generate flows, on the order of 10 seconds or greater. Second, our times are subjective, with a participant-reported indication of being “done.” Third, post-interview data indicates that some participants felt they could go further in their analysis with the assistant, compared to without—for instance, whereas 6 participants ran evaluators and spent time inspecting evaluation results with the assistant, only a single participant without the assistant ran evaluators. Finally, though participants spent around the same time per condition, what they spent that time on differed. Participants using the assistant spent time clarifying their intent with the assistant, revising, running or inspecting the generated flows. Interaction traces (Fig. 6) of user actions indicated that considerably less time was spent on preparing and editing input data for the pipeline, in particular.
More broadly than LLM pipelines, our work contributes to a growing body of literature on AI agents integrated into software platforms to assist users in implementing ideas (e.g., an agent that creates Powerpoint slides, visualizations from datasets, or dynamic widgets (Vaithilingam et al., 2024b; Beasley and Abouzied, 2024; Zheng et al., 2022; Zhou et al., 2023; Cheng et al., 2024)). Our study provides evidence for the intuitive hypothesis that users perceive that AI agents improve workload by reducing the mental and physical burdens placed on them when interacting with software—clicking and dragging, planning what to do. However, our quantitative findings indicate that for the most part, users perceived they were just as successful without the assistant—even when performance analysis suggests performance plummeted without the assistants’ help. This finding resembles the Dunning–Kruger effect (Kruger and Dunning, 1999)—that people who lack the skill to evaluate the quality of their work in a domain may over-estimate its quality—and, combined with worries about being biased by the assistant (discussed below), contributes to growing concerns about non-experts’ over-reliance on AI systems (Buçinca et al., 2021; Bansal et al., 2021). Note that prior studies on LLM-assisted workflow generation only report user self-perceptions of performance (e.g., (Zhou et al., 2023; Cheng et al., 2024)); our findings suggest future studies on workflow generation assistants need to complement subjective measures with more objective evaluations of performance (e.g., by expert raters).
@@ 6.1. Risk of over-reliance on AI assistance for LLM pipeline creation
The reader might ask whether and to what extent the user over-relies on, or is biased by, the assistant’s solutions. Even if the assistant helps the user learn the platform (which some participants indicated), is there a risk of it proposing solutions that are subpar or misleading in some way, outside of more straightforward usability errors? For instance, imagine a user asks ChainBuddy to set up a flow to evaluate LLMs for gender bias. In this task, ChainBuddy usually sets up a flow that compares whether the presence of a social identity marker (male, female, non-binary) changes the LLM success rate on solving a math problem—a common “persona” approach used in benchmarking papers (Cheng et al., 2023). While one way to evaluate gender bias, it certainly does not exhaust all possibilities, nor does the current assistant help the user “scale up” their analysis, choose one solution from alternatives and weigh trade-offs (a learning technique from variation theory (Ling Lo, 2012)), or help them hedge against misinterpretation of random noise. In our study, we did see one error case where ChainBuddy constructed a wrong solution that did not compare two prompts. The participant was evidently not able to identify and fix the AI’s mistakes. This raises the question of how future assistants can help the user learn the interface, while mitigating the potential for bias.
6.2. What is the right interaction design for gathering user requirements?
AI-assisted workflow generation is a budding area of research that requires eliciting user requirements. However, it remains unclear what the relative trade-offs are for different interface designs that elicit user intent. Past approaches allow only for a single prompt (Zhou et al., 2023; Kulkarni et al., 2012), adopt an “infinite questioner” that ask one question at a time until the user presses generate (Cheng et al., 2024), or stop generation in stages to ask for user feedback before continuing (Kazemitabaar et al., 2024). Here, we explored a structured form-filling approach that mixed chat with a three-question form, where users could choose which questions to answer; in part, this was because our early pilot studies suggested that an unstructured infinite questioner was time-consuming and pressuring to users. Users in our study appreciated our approach, however it still remains unclear what the “right” approach is to eliciting user requirements. In particular, we felt that while functional, the form-filling approach feels awkward inside a chat window. Future work could investigate the advantages and drawbacks of different interfaces for eliciting user requirements for generation tasks with high complexity, and perhaps compare findings with interactive “steering”-based approaches such as Kazemitabaar et al. (Kazemitabaar et al., 2024).
@@ 6.3. Future Work
Appendix A Technical Evaluation Details
Here we report more details and results from our a technical evaluation (Section 3.4).
A.1. Expert Rating Guidelines
Here is the prompt we gave the expert raters for their task:
“Score the workflow quality from 1 to 5, with 5 the highest score and 1 the lowest score, based on how well the workflow satisfied the requirement(s) in the task description. Keep in mind that these workflows are meant to pose a good starting point to kickstart user’s explorations. Also focus on the quality of the workflow, rather than the quality of the interface.”
To help align experts in the under-specified domain of workflow evaluation, we also provided the following specific descriptions of what constitutes a score of 1 (lowest and least aligned to the user requirements) to 5 (most aligned with no errors and no fixes required):
  1. (1)
    Not aligned to the requirements at all.
  2. (2)
    Slightly aligned to the requirements, but with some significant misunderstanding, missing features, or error.
  3. (3)
    Aligned to the requirements. The workflow is a good starting point for users, but has at least one significant problem or missing feature (either missing an important part, or having 1-2 incorrect implementations of nodes or connections).
  4. (4)
    Aligned to the requirements, with only 1-2 minor issues. Any errors are easily fixable and the workflow is extensible.
  5. (5)
    Strongly aligned to the requirements. The workflow represents a great starting point for the user, and does not require any fixes to its existing content.
A.2. Results from Technical Evaluation
In the table below, we report all prompts from our technical evaluation, alongside rater scores and averaged scores, with the qualitative feedback summarized in the Issues column.
Workflow Prompt S1 S2 Avg.Score Summarized Issues
Prompt Engineering - Compare a prompt across models
Q1.1.1 Design a workflow to test a single prompt across GPT-4, GPT-3.5, and Claude for creative writing tasks. Visualize response length and diversity. 5 5 5.0 No outstanding issues.
Q1.1.2 ”I want to compare how GPT-4 and Claude respond to a prompt that provides recipes for making different types of pies. Then, I want to easily evaluate the number of steps in each recipe.” 3 5 4.0 Included a prompt comparison when it should not.
Q1.1.3 ”I want to see how different models respond to the same prompt for creating an imaginary conversation between characters in a video game.” 5 3 4.0 One rater felt evaluation is extensive and appropriate, but another thought it was too extensive (”feels bloated”) with too little inputs and too many evaluators.
Prompt Engineering - Compare different prompts
Q1.2.1 I want to compare different prompts for understanding how professionals from different backgrounds overcome challenges in their lives. 4 4 4.0 Minor issues to Python evaluator, Vis Node choice of variable to plot; tweak needed to prompt in Prompt Node.
Q1.2.2 ”Make a workflow to compare different prompts for shortening text while keeping the same content. Check and compare the length of the text across prompts.” 5 5 5.0 Input data (example paragraphs) could be longer to test.
Q1.2.3 ”Set up a workflow to compare prompts for generating formal yet concise emails given user descriptions of what the email should be about.” 4 3 3.5 Could make minor improvement to prompt to clarify expected structure of output; LLM evaluator prompt ”seems off” based on the high scores and might need further alignment
Prompt Engineering - Test the robustness of prompt across different inputs
Q1.3.1 Can you help me check how my prompt responds to many different inputs? The prompt is: “Give me a fun recipe for dessert.” I’d like to try this prompt across different desserts. 5 4 4.5 Python evaluator’s objective is unclear.
Q1.3.2 Generate a workflow evaluating how commands like ’Be formal’ or ’Be casual’ influence conversational model outputs. Include a plot of response sentiment. 4 5 4.5 Visualization should plot on a different variable.
Q1.3.3 I have a prompt for rephrasing verbose sentences and want to test for robustness across many different inputs. Here is the prompt: “Rephrase the following sentence to be more concise but keep the same meaning: sentence.” 5 2 3.5 Missing an evaluator node to check for length; one rater felt there was a structural issue with the template chaining setup
Table 1. Prompt Engineering Evaluation
Workflow Prompt S1 S2 Avg.Score Summarized Issues
Auditing - Checking accuracy
Q2.1.1 Create a workflow to benchmark LLMs using a set of trivia questions. Evaluate responses for correctness. 5 3 4.0 Python evaluator logic must be amended.
Q2.1.2 ”I want to compare how different models reply to the same math question across three different languages: English, German, and Russian.” Evaluate the accuracy of the response.” 4 5 4.5 Evaluation is limited only to Russian; evaluator choice could be better
Q2.1.3 ”Design a workflow to test factual accuracy of LLM responses on historical questions about Ancient Rome.” 5 4 4.5 One rater felt that LLM scorer was suboptimal and instead the system should use a RAG approach [note: the baseline interface does not have support for evaluating RAG]
Auditing - Bias Detection
Q2.2.1 Generate a workflow to detect cultural bias in LLM outputs by varying names and locations in prompts. 5 5 5.0 LLM scorer prompt should be refined.
Q2.2.2 “I want to check whether giving a model a persona across different gender and racial identities changes how well they reply to mathematical problems.” 3 4 3.5 Includes an unnecessary LLM scorer not asked for by the user; visualization node should plot a different variable
Q2.2.3 “Have the model describe different jobs by only varying demographic variables (e.g., gender, ethnicity) and see if responses differ across these variables. For instance, you might compare how it describes a female doctor’s job versus a male doctor. Grade the sentiment of the outputs.” 4 3 3.5 Prompt in LLM scorer needs to be revised; input data could vary more demographic variables
Auditing - Ethics and Red-teaming
Q2.3.1 I want to check how different models respond to ethical dilemmas and whether they make more altruistic choices or not. Make it easy to evaluate the altruism of the response 5 5 5.0 No outstanding issues.
Q2.3.2 “I’m curious in what circumstances different models will refuse to respond given their safety filters. Generate a workflow to help me check across models whether the model refuses to reply to a given question. Evaluate whether the output is a refusal to answer the question or not.” 5 4 4.5 Evaluation could be improved by adding more types of evaluations; visualization node should plot a different variable
Q2.3.3 “Design a workflow to help me classify the toxicity of social media comments, across different possible prompts to GPT-4. Score the toxicity on a scale of 1-5, with 5 the most toxic, and plot the results so that I can compare.” 3 4 3.5 Prompts vary in quality; could possibly add multiple models.
Table 2. LLM Auditing Evaluation
Workflow Prompt S1 S2 Avg.Score Summarized Issues
Data Processing - Process and clean input data
Q3.1.1 Build a workflow to clean up ASR transcript excerpts that have some mistakes, revising the transcripts to correct grammar or spelling issues but not changing the meanings 4 5 4.5 Missing evaluation to double-check the output of the LLM.
Q3.1.2 “I want to extract information from different emails and output a JSON format with dictionary keys “importance”, “sender”, “recipient”, and “description”. 5 4 4.5 Should have use a Python evaluator instead of the LLM Scorer to check JSON format of outputs; prompt could be clarified slightly in what the definition of ”importance” is
Q3.1.3 “I’m curious if asking for different formatting (e.g., asking a model to reply in JSON, YAML, or Markdown) impacts the content of the outputs. Make a workflow to help explore this.” 4 4 4.0 Visualization should plot on different variable; LLM scorer prompt needs refinement
Data Processing - Data Augmentation and Generation
Q3.2.1 Give me a workflow that generates paraphrased versions of sentences for data augmentation in sentiment analysis. 5 5 5.0 No outstanding issues.
Q3.2.2 ”I want to use LLMs to generate example long paragraphs for a summarization task. Use multiple models for some diversity.” 4 3 3.5 Could have added more models; might have added an LLM scorer to double-check quality of outputs
Q3.2.3 “Design a workflow that converts all input questions from English into a target language (e.g., French, Korean). Evaluate the language of the response to double-check.” 4 4 4.0 Visualization should plot on different variable; could have explored more prompt variations for the translation task
Data Processing - Entity Extraction
Q3.3.1 Create a workflow to extract all dates and locations mentioned in LLM responses to a historical question. 2 2 2.0 The prompt template in Prompt Node is not aligned; and the evaluation is incorrect/an approach that will struggle to generalize
Q3.3.2 “I want to check how well different models understand different famous inventors and the entities that they put in their responses. Extract the named entities so that I can compare.” 3 3 3.0 Does not compare different models in the Prompt Node, which was a necessary requirement, but otherwise appropriate
Q3.3.3 “I want to extract the names and public figures mentioned in some sample tweets. Exclude entities which appear in hashtags and output the entities as a comma-separated list.” 3 5 4.0 Code evaluator is ”too stringent”, does not address enough edge cases; visualization node plot appears empty
Table 3. Data Processing Evaluation
Appendix B Walkthrough of System Architecture: Comparing Prompts for Generating Catchy Tweets
Here we walk through an example of how our system architecture works, from the initial user requirements (prompt) to the generation of a workflow.
Imagine a user provides the following input to the system, which is similar to Task 2 in our user study:
”I want to compare different prompts that can generate catchy tweets from long paragraphs. The generated tweets should be concise, staying within the 144-character limit.”
The Planner Agent receives the user requirement as context to a long prompt (Figure 7). The Planner prompt decomposes the workflow into suggested nodes segmented by layers (described below), assigns variables, and traces connections across nodes. First, it generates the workflow skeleton following a layer-based design that segments the parts of evaluative LLM pipelines into three general sections (B.2):
  • Input Layer: Provides fixed inputs to the workflow that do not include any template variables.
  • Middle Layer(s): Process and transform input data using template variables and (potentially) prompt chaining. Includes nodes that generate prompts, process responses, or chain variables together.
  • Evaluation Layer: Includes evaluator nodes for assessing outputs generated by the workflow.
These layers occur in sequence (left-to-right) in the resulting node-edge graph. The reason we use this layer-based design is because later layers like Evaluation depend upon the results of earlier layers (e.g., the type of input data and the specific names of prompt variables). More detail on how we described the layers in the system prompt can be found in Section B.2.
[…]
This figure is a visual representation of a planner agent system’s prompt, designed with a structured format and color-coded text for clarity. The header at the top is labeled ”Planner” on a light yellow background. Below, the content is divided into sections: ###Role: You are a Planner. Your responsibility is to create a detailed plan to help your team set up a correct chain …… ###Workflow Design Principles: Think in Layers: [Explanation of how the workflow is divided into Input, Middle, and Evaluation Layers.] Node Responsibilities: [Available Nodes in each layer.] [Usage of TextFieldsNodes.] [Usage of Intermediate PromptNodes.] Variable Traces: [Explanation of how variables are tracked across the workflow.] Reasoning and Skeleton Design: [Details of how the Planner creates a skeleton workflow design to justify the structure.] ###Process Understand Requirements: [Description of analyzing user requirements.] Layer-Based Design: [Steps to divide the workflow into logical layers and design connections.] Justify Design: [Description of how to validate the workflow against user objectives.] Output the Plan: [Steps to generate the final JSON plan, including nodes, variable traces.] ###examples:[example]
Figure 7. Planner agent system prompt. Note that this prompt is very long and hence truncated for length here, with details for each section summarized in square braces.
For instance, here is example JSON output of the Planner outlining its high-level workflow plan to address the user requirement via four layers (note that these example outputs are condensed for length):
"layer_skeleton": {
"justification":
"This workflow transforms long paragraphs into catchy tweets.",
"design": [
{ "layer_id": 1, "name": "Input Layer", "description": "Collects long paragraphs for tweet generation." },
{ "layer_id": 2, "name": "Template Layer", "description": "Generates prompt templates for creating concise tweets." },
{ "layer_id": 3, "name": "Prompt Node Layer", "description": "Queries LLMs using generated prompts to create tweets." },
{ "layer_id": 4, "name": "Evaluation Layer", "description": "Evaluates the tweets for adherence to character limits." }
Expanding on this high-level plan, the Planner then provides implementation examples of nodes corresponding to each layer for the agents to get started. Task-specific agents are then provided with both the high-level plan and each task-specific example that relates to their task, but are not required to stick to the example. The implementation examples is an array of JSON objects like the following:
"layer_id": 1,
"name": "Input Layer",
"nodes": [
"node_id": "TextFieldsNode1",
"type": "TextFieldsNode",
"description": "Define long paragraphs as input",
"variables": {
"long_paragraph": [
"The quick brown fox jumps over the lazy dog...",
"In a world where technology is advancing rapidly...",
"The art of communication is not just about speaking..."
"layer_id": 2,
"name": "Template Layer",
"nodes": [
"node_id": "TextFieldsNode2",
"type": "TextFieldsNode",
"description": "Create prompts for generating catchy tweets",
"variables": {
"prompt_template_tweet": [
"Summarize the following into a catchy tweet: {long_paragraph}",
"Create a tweet from this paragraph, keeping it under 144 characters: {long_paragraph}"
... ### Other layers
Finally, the Planner agent generates a list of template variable names used in the flow, what we call variable tracing. Variable tracing ensures each variable is clearly defined, connected, and traceable throughout the workflow, which is particularly necessary for template chaining. Moreover, the tracing allows agents to identify which nodes generate or consume specific variables, making it easier to validate the workflow and reuse any previous information during evaluation. The example variable trace produced by the Planner is an array of JSON objects, e.g.:
"variable": "{long_paragraph}",
"origin": "TextFieldsNode1",
"consumed_by": ["TextFieldsNode2"]
"variable": "{prompt_template_tweet}",
"origin": "TextFieldsNode2",
"consumed_by": ["PromptNode1"]
"variable": "{tweet_response}",
"origin": "PromptNode1",
"consumed_by": ["EvaluatorNode1", "LLMScorerNode1"]
After the planner generates its plan, its output is stored in a global state. Each task-specific agent then gets their task list by using tools to retrieve the corresponding plan for each node of their task type. Here are broken-down examples for each agent after the Planner generates the plan for the example user requirements:
B.0.1. TextFields Agent
This agent receives a task list that contains all TextFieldsNode specifications from the Planner’s output, and populates TextFieldsNode with input data based on the task list and ensures correct braces are used for specific variables.
[…]
The figure shows the workflow of given a task list to the TextFields agent and elaborating it into a complete specification in JSON format. The layout is structured as follows: The top section contains two labeled boxes connected by a blue arrow: Task list (on the left, light yellow background): Represents the Planner agent’s output, detailing tasks to be implemented: ”layers”: [ ”layer_id”: 1, ”name”: ”Input Layer”, ”nodes”: [ ”node_id”: ”TextFieldsNode1”, ”type”: ”TextFieldsNode”, ”description”: ”Define long paragraphs as input”, ”variables”: ”long_paragraph”: [ ”The quick brown fox jumps over the lazy dog……”, ”In a world where technology is advancing rapidly……”, ”The art of communication is not just about speaking……” ] ] ,…###Remain tasks textAgent (on the right, light yellow background, with a laptop emoji): Represents the TextFields agent, which receives the task list and processes it with output: ”Nodes”: [ ”id”: ”TextFieldsNode1”, ”data”: ”fields”: ”f1”: ”The quick brown fox jumps over the lazy dog. This sentence is a pangram, meaning it contains every letter of the alphabet.”, ”f2”: ”In a world where technology is advancing rapidly, staying updated with the latest trends is crucial for success.”, ”f3”: ”The art of communication is not just about speaking but also about listening and understanding the other person’s perspective.” , ”title”: ”Long Paragraphs” , … ###Remain output Remain tasks (gray text at the bottom left): Placeholder for any remaining tasks to be implemented by the TextFields agent. Remain output (gray text at the bottom right): Placeholder for remain specifications generated by the TextFields agent.
Figure 8. TextFields agent takes the task list from the Planner agent with its example sketch of the node in JSON, and implements it with a complete specification.
B.0.2. Prompt Agent
This agent receives a task list containing all PromptNode specifications from the Planner’s output, and configures PromptNodes to query multiple LLMs (GPT-4, Claude, etc.).
[…]
The figure shows the workflow of given a task list to the Prompt Agent and elaborating it into a complete specification in JSON format: The layout is structured as follows: The top section contains two labeled boxes connected by a blue arrow: Task list (on the left, light yellow background): Represents the Planner agent’s output, detailing tasks to be implemented
Figure 9. Prompt agent takes the task list from the Planner agent, and implements Prompt Nodes with a complete specification.
B.0.3. Evaluator Agent
This agent receives user requirements, example prompts provided from the previous agents, and potential responses as examples of LLM outputs that will be fed into evaluator:
[…]
This figure displays the structured system prompt for the ”Evaluator Agent.” The prompt uses a clear format to delineate sections and emphasize key processes and rules. The prompt is also truncated for better display. The header is labeled Evaluator Agent on a light yellow background.
Figure 10. Evaluator agent system prompt (truncated for length)
The evaluator agent then produces one or more evaluation nodes that can be a mix of Python Code Evaluators and LLM Scorers, expressed as JSON objects:
[…]
Figure 11. Evaluator agent input and output for the example. For the input (left), we provide example prompts and LLM outputs in addition to the user requirements, to help the evaluator to anticipate the general format of LLM responses that will be passed into it. This agent produces two evaluators: one Python Code Evaluator and one LLM Scorer.
B.0.4. Edge Agent
This agent receives variable traces from the Planner’s output, and ensures proper data flow across the workflow by establishing edges between nodes in the format expected by ChainForge:
[…]
This figure visualizes the input and output of the EdgeAgent, which is responsible for creating edges in the JSON format. Input (blue arrow pointing to the right): Represents the data sources for the EdgeAgent: Variable traces from the Planner agent. EdgeAgent (right side, light yellow background with a laptop emoji): Represents the computational agent that processes the input and constructs a list of edges. These edges define connections between the nodes and are formatted in the JSON structure expected by the ChainForge system. ”edges”: [ ”source”: ”TextFieldsNode1”, ”sourceHandle”: ”output”, ”target”: ”TextFieldsNode2”, ”targetHandle”: ”long_paragraph”, ”id”: ”…” , ”source”: ”TextFieldsNode2”, ”sourceHandle”: ”output”, ”target”: ”PromptNode1”, ”targetHandle”: ”prompt_template_tweet”, ”id”: ”…” , ”source”: ”PromptNode1”, ”sourceHandle”: ”prompt”, ”target”: ”EvaluatorNode1”, ”targetHandle”: ”tweet_response”, ”id”: ”…” ]
Figure 12. Edge agent takes the variable traces from the Planner agent and the generated nodes, and constructs a list of edges in the JSON format expected by ChainForge.
B.0.5. Layout Agent
This agent receives output from Edge Agent and its responsibility is to simply configure every nodes’ position coordinates (from left to right) for the final workflow.
B.1. Final Workflow
Once all agents have completed their tasks, the workflow is assembled with a default configuration template for each type of node and then merged into a single JSON file which is then presented in the interface (see Fig. 2 for an example generation that addresses the same tweet task as here).
B.2. Layer-Based Workflow Skeleton
To generate the entire workflow, the Planner agent translates user requirements into a structured, layer-based workflow skeleton. Here we provide further details of how we describe layers in the Planner’s system prompt (Fig. 7).
Layer Definitions:
Input Layer (Layer 0):
  • Purpose: Provides fixed inputs to the workflow that do not include any template variables.
  • Structure: Consists of nodes (e.g., TextFieldsNode) that define raw, static inputs. These inputs serve as the foundation for generating prompts in subsequent layers.
  • Characteristics:
    • Contains data directly entered by the user (e.g., email content, comments, or datasets).
    • No dynamic variables are used within this layer.
Middle Layers (Layer 1, Layer 2, etc.):
  • Purpose: Process and transform input data using template variables and prompt chaining.
  • Structure: Includes nodes that generate prompts, process responses, or chain variables together.
  • Characteristics:
    • Dynamically incorporates variables generated in earlier layers.
    • Variables and nodes are connected in a sequential, logical flow to ensure proper data transformation.
    • Each variable has a unique ID to maintain clarity and prevent conflicts across layers and nodes.
Evaluation Layer (Final Layer):
  • Purpose: Includes evaluator nodes for assessing outputs generated by the workflow.
  • Structure: Contains nodes (e.g., LLMScorerNode, CodeEvaluatorNode) designed to measure outputs against predefined criteria.
  • Characteristics:
    • Designs different numbers and types of evaluators based on below context given to the evaluator agent:
        User Requirements: Direct criteria for evaluation (e.g., ”rate professionalism”).
        Example Prompts: Example prompts inside current workflow.
        Potential Responses: Context-aware example LLM outputs that evaluators may expect.
Appendix C Participant Ideas in Free Exploration